hi!!!!! this is my first year in algebra and i have a question.... my question is " evaluate 1/3 (10-7) +4 (18 divided by 9)
then it has problems on the bottom it says name the property in each step
1/4(10-7)+4(18dividedby9-1)
=1/3(3)+4(18 divided by 9 - 1)____________
=1/3 (3) + 4 (2-1)___________
=1/3 (3) + 4(1)_____________
=1+ 4____________
=5____________
: hi!!!!! this is my first year in algebra and i have a question.... my question is " evaluate 1/3 (10-7) +4 (18 divided by 9)
then it has problems on the bottom it says name the property in each step
1/4(10-7)+4(18dividedby9-1)
=1/3(3)+4(18 divided by 9 - 1)____________
=1/3 (3) + 4 (2-1)___________
=1/3 (3) + 4(1)_____________
=1+ 4____________
=5____________

Answer by stanbon(19020)

evaluate 1/3 (10-7) +4 (18 divided by 9)
= (1/3)(3) + 4(2)
= 1 + 8
= 9
-------------------------
then it has problems on the bottom it says name the property in each step
1/3(10-7)+4(18dividedby9-1)
=1/3(3)+4(18 divided by 9 - 1) evaluate a parenthesis____________
=1/3 (3) + 4 (2-1) simplify a parentheses___________
=1/3 (3) + 4(1) simplify a parentheses_____________
= 1 + 4 simplify ____________
=5 simplify ____________
-----------------------------
Comment: Usually these problems involve commutative, distributive,
associative, and identity properties. Your problem seems to involve
only arithmetic simplification.
==============================
Cheers,
Stan H.